OpenStudy (anonymous):

Help? After a semester you complete 20 homework assignments with an average of 17. When your lowest score of 12 is dropped, what is your new average?

OpenStudy (dumbcow):

avg = sum/N ---> sum = N*avg sum = 20*17 = 340 score of 12 is dropped 340 -12 = 328 avg = 328/16